OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 4. Find x when y = 8.

OpenStudy (bahrom7893):

y varies inversely as 3 times x means: 3 times x = 3*x inversely as 3*x = 1/(3x), so y = 1/(3x)

OpenStudy (bahrom7893):

Well actually y = a/(3x), we don't know what a is yet.

OpenStudy (bahrom7893):

But we do know that when x = 6 y=4 4 = a/(3*6) => a = 4*3*6 = 72

OpenStudy (bahrom7893):

y = 72/(3x) = 24/x

OpenStudy (bahrom7893):

when y = 8: 8 = 24/x x = 3

OpenStudy (bahrom7893):

I think that's right, satellite, can you check it over please?

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

looks good to me

OpenStudy (bahrom7893):

tnx